OpenStudy (anonymous):

Factor: x2 - 11x + 18 (x + 1)(x - 12) (x - 3)(x - 6) (x - 2)(x - 9) prime

OpenStudy (anonymous):

(x-9)(x-2)

OpenStudy (hoblos):

x2 - 11x + 18 = x2 - 2x - 9x +18 = x(x-2) -9(x-2) = (x-2)(x-9)